OpenStudy (18jonea):

Describe how to transform the graph of f(x) = logx into g(x )=3 logx. Stretch the graph of f(x) by a factor of 3. Shift the graph of f(x) right 3 units. Shrink the graph of f(x) by a factor of 3. Shift the graph of f(x) up 3 units. @3mar I think a

OpenStudy (3mar):

Plot them and notice!

OpenStudy (18jonea):

so im almost positive it is a

OpenStudy (3mar):

I think this would help you.

OpenStudy (18jonea):

well you dont move the graph the right or up 3 units so b and d are out

OpenStudy (18jonea):

right @3mar

OpenStudy (3mar):

1 min

OpenStudy (18jonea):

k

OpenStudy (steve816):

@3mar the picture you posted doesn't have anything about shrink or stretch.

OpenStudy (3mar):

so the graph is!

OpenStudy (18jonea):

i still think a is a incorrect?

OpenStudy (3mar):

For Stretch the graph of f(x) by a factor of 3................=3log(x) For Shift the graph of f(x) right 3 units.........................=log(x-3) For Shrink the graph of f(x) by a factor of 3..................=\(\frac{1}{3}\)log(x) For Shift the graph of f(x) up 3 units.............................=log(x)+3 I hope I helped you.

OpenStudy (18jonea):

yes so the answer is a

OpenStudy (18jonea):

@3mar

OpenStudy (3mar):

\[\Huge\color{red}\checkmark\]

OpenStudy (3mar):

Thank you for the medal! Are you satisfied?

OpenStudy (18jonea):

i have one more. yes very satisfied

OpenStudy (3mar):

happy to hear that Go on I am with you!